OpenStudy (anonymous):

quadratic equation: square root property: 6x^2-24=0

OpenStudy (anonymous):

6x^2 - 24 = 0 6x^2 = 24 x^2 = 4 x = +/- 2

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

your welcome:]

OpenStudy (anonymous):

how about? 3x^2 - 8 = 4

OpenStudy (anonymous):

i dont know how to solve this 3x^2 - 8 = 4

OpenStudy (anonymous):

x=+/- 2 ?? am i right?

OpenStudy (anonymous):

Yeah i think soo:]